OpenStudy (anonymous):

A sporting goods store has added a section with trail foods and snacks. Raisins cost $5.00 per pound and peanuts cost $3.80 per pound. Raisins and peanuts are going to be sold in a mixture of 30 pounds which will cost $4.05 per pound? How many pounds of raisins will be in the mix

OpenStudy (kropot72):

r = number of pounds of raisins p = number of pounds of peanuts r + p = 30 (5*r) + (3.8*p) = 4.05 * 30 = 121.5 Multiply both sides of the equation r + p = 30 by 5: (5*r) + (5*p) = 150 and subtract corresponding terms of (5*r) + (3.8*p) = 121.5 from it giving: 1.2 * p = 28.5 p = 28.5/1.2 = 23.75 pounds of peanuts r = 30 - 23.75 = 6.25 pounds of raisins